OpenStudy (anonymous):

Can someone please help me find an explicit formula for the sequence 1,2,4,7,11,16,22... ?

OpenStudy (tkhunny):

Subtract each value after the first from the number that preceeds it. 2 - 1 = ?? 4 - 2 = ?? 7 - 4 = ?? Do them all and see if you notice a pattern.

OpenStudy (anonymous):

Yurps Nice job didnt get it at 1st

OpenStudy (anonymous):

I have already got that far. I know the pattern increases by one more than the first each time. I have even made the recursive formula for it (I think). \[a _{n}= a _{n-1} + n\]But I do not know how to make the explicit formula. I am repeatedly getting stuck.

OpenStudy (anonymous):

Wait, no. I don't think that formula is right either...

OpenStudy (anonymous):

Is something wrong with my sequence? is it even formula-able? O.o

OpenStudy (anonymous):

I'll check back... Hopefully someone can help me!

OpenStudy (tkhunny):

You must learn DISCOVERY. Try some things. \(a_{n} = n^{2}\) 1 4 9 16 ... Okay, that's a little too fast. \(a_{n} = n^{2} - n\) 0 2 6 12 ... Off by at least 1 at the beginning \(a_{n} = n^{2} - n + 1\) 1 3 7 13 ... I see it. \(a_{n} = ½(n^{2} - n + 2)\) It takes a little effort. Note: that little experiment at the beginning told us that a quadratic function would be sufficient. We didn't just do that for fun.

OpenStudy (anonymous):

Thank you, truly. That really helped me understand. :)